OpenStudy (anonymous):

Which of the following is a solution of x2 - 6x = -10? x = -3 - i x = 3 + i x equals 3 plus i times the square root of 19 x equals negative 3 minus i times the square root of 19

OpenStudy (charlotte123):

x = 3 + i ^_^ http://www.wolframalpha.com/input/?i=x%5E2+-+6x+%3D+-10 :D

OpenStudy (solomonzelman):

\[x2 - 6x = -10\]\[x2 - 6x + 10 = -10 +10\]\[x^2 - 6x +10=0\] Which methods to solve this do you know?

OpenStudy (solomonzelman):

Factoring Quadratic formula Completing the square

OpenStudy (solomonzelman):

\[x^2 - 6x+ 10=0\]\[(x-3)^2=-1\]technically\[x=±i+3\]so\[x=i+3\]